# Rethinking Approximate Gaussian Inference in Classification

### v3

@@ E.1.1 Softmax
The Mean Field predictive [lu_mean-field_2021] uses the following approximation for the true predictive:
@@
The Laplace Bridge predictive [pmlr-v180-hobbhahn22a] approximates the true predictive as follows:
@@
The shekhovtsov_feed-forward_2018 approximation of the Gaussian-softmax integral yields:
@@ E.2.1 Softmax
The Laplace Bridge method [pmlr-v180-hobbhahn22a] uses the following Dirichlet parameters:
@@ Appendix G Experimental Setup
To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [pmlr-v180-hobbhahn22a], Mean Field [lu_mean-field_2021], Monte Carlo sampling). We consider Heteroscedastic Classifiers (HET) [collier2021correlated], Spectral-Normalised Gaussian Processes (SNGP) [liu2020simple], and last-layer Laplace approximation methods [daxberger2021laplace] as backbones (see Appendix˜J for details).
The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [deng2009imagenet] and CIFAR-10 [krizhevsky2009learning] on five metrics aligning with practical needs from uncertainty estimates [mucsanyi2023trustworthy]:
@@
For ImageNet, we treat ImageNet-C [hendrycks2019benchmarking] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
@@
For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking. We train with the LAMB optimiser [you2019large] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following tran2022plex. We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$. The learning rate is treated as a hyperparameter and selected from the interval $[0.0005,0.05]$ based on the validation performance. The weight decay is selected from the set $\{0.01,0.02\}$. During training, we keep track of the best-performing checkpoint on the validation set and load it before testing. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [shahriari2015taking] in Weights & Biases [wandb] based on the negative log-likelihood.
On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The only exceptions are the SNGP models that are trained for 125 epochs [liu2020simple]. We train with Momentum SGD using a batch size of $128$ and no gradient accumulation. Similarly to ImageNet, we use a cosine learning rate schedule but with five warmup epochs and warmup learning rate $1\mathrm{e}{-5}$. The learning rate is also treated as a hyperparameter on CIFAR-10. We use the interval $[0.05,1]$ for Sigmoid and NormCDF, and $[0.01,0.15]$ for Softmax and Exp. The optimal learning rates for Sigmoid and NormCDF are generally larger, as the class-wise binary cross-entropies are averaged instead of summed. The weight decay is selected from the interval $1\mathrm{e}{-6},1\mathrm{e}{-4}]$. Similarly to ImageNet, we use the best-performing checkpoint in the tests and use a ten-step Bayesian Optimization scheme to select performant hyperparameters.
@@ I.1 Log Probability Proper Scoring Rule for the Predictive
First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [mucsanyi2023trustworthy].
@@ I.2 Expected Calibration Error
To set up the required quantities for the Expected Calibration Error (ECE) metric [naeini2015obtaining], we follow the steps below, based on [mucsanyi2023trustworthy].
@@
The Expected Calibration Error (ECE) metric [naeini2015obtaining] is then defined as
@@ I.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
One can show that this is indeed a strictly proper scoring rule [mucsanyi2023trustworthy].
@@ I.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
Spectral normalised Gaussian processes (SNGP) [liu2020simple] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty. For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
@@
The method applies spectral normalization to the hidden weights in each layer using a power iteration scheme with a single iteration per batch to obtain the largest singular value. liu2020simple claim this helps with input distance awareness.
@@
Heteroscedastic classifiers (HET) [collier2021correlated] construct a Gaussian distribution in the logit space to model per-input uncertainties:
@@
The Laplace approximation [daxberger2021laplace] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian. In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularised loss over the training set $\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
@@
This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [tatzel2025debiasing].
For deep neural networks, the Hessian matrix is often replaced with the Generalised Gauss-Newton (GGN) matrix. The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [martens2015optimizing] or low-rank approximations.
@@
We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [tatzel2025debiasing].
@@
Table˜K.3 shows results on ViT Little [dosovitskiy2020image] backbones from the timm [rw2019timm] library on the ImageNet validation set. As in the main paper, the hyperparameters are optimised using a ten-step Bayesian hyperparameter optimization scheme of Weights and Biases.
@@
We first evaluate calibration using the log probability scoring rule [gneiting2007strictly] and the Expected Calibration Error (ECE) metric [naeini2015obtaining]. Fig.˜3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
@@
Closed-form predictives do not sacrifice accuracy. Fig.˜K.3 evidences this claim on CIFAR-10: our closed-form predictives either outperform or are on par with Softmax predictives. The most accurate method is Sigmoid HET. These results support the findings of wightman_resnet_2021 that showcase desirable training dynamics of the class-wise cross-entropy loss.
@@
Table˜K.8 provides results for two evidential deep learning methods, EDL [sensoy2018evidential] and PostNet [charpentier2020posterior]. (We adopt the naming convention of [mucsányi2024benchmarking].) We make the comparison because these methods also yield closed-form predictives and second-order quantities of interest, as they build Dirichlet distributions whose advantages we discuss in Section˜4.1. Following our experimental setup (Section 6), we optimise the hyperparameters of both methods using a ten-step Bayesian Optimization scheme in Weights & Biases. Compared with the results presented in the main text, our approach outperforms the two EDL methods across all metrics, significantly so on NLL and OOD AUROC. The results on the latter metric indicate that our moment-matched Dirichlet distributions capture second-order epistemic information more effectively than the loss-based EDL methods, which aligns with the findings of NEURIPS2022_bc1d640f.

### v4

@@ E.1.1 Softmax
The Mean Field predictive [31] uses the following approximation for the true predictive:
@@
The Laplace Bridge predictive [19] approximates the true predictive as follows:
@@
The Shekhovtsov and Flach [46] approximation of the Gaussian-softmax integral yields:
@@ E.2.1 Softmax
The Laplace Bridge method [19] uses the following Dirichlet parameters:
@@ Appendix G Experimental Setup
To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [19], Mean Field [31], Monte Carlo sampling). We consider Heteroscedastic Classifiers (HET) [7], Spectral-Normalised Gaussian Processes (SNGP) [30], and last-layer Laplace approximation methods [11] as backbones (see Appendix˜J for details).
The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [12] and CIFAR-10 [25] on five metrics aligning with practical needs from uncertainty estimates [38]:
@@
For ImageNet, we treat ImageNet-C [18] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
@@
For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking. We train with the LAMB optimiser [55] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following Tran et al. [50]. We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$. The learning rate is treated as a hyperparameter and selected from the interval $[0.0005,0.05]$ based on the validation performance. The weight decay is selected from the set $\{0.01,0.02\}$. During training, we keep track of the best-performing checkpoint on the validation set and load it before testing. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [45] in Weights & Biases [5] based on the negative log-likelihood.
On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The only exceptions are the SNGP models that are trained for 125 epochs [30]. We train with Momentum SGD using a batch size of $128$ and no gradient accumulation. Similarly to ImageNet, we use a cosine learning rate schedule but with five warmup epochs and warmup learning rate $1\mathrm{e}{-5}$. The learning rate is also treated as a hyperparameter on CIFAR-10. We use the interval $[0.05,1]$ for Sigmoid and NormCDF, and $[0.01,0.15]$ for Softmax and Exp. The optimal learning rates for Sigmoid and NormCDF are generally larger, as the class-wise binary cross-entropies are averaged instead of summed. The weight decay is selected from the interval $1\mathrm{e}{-6},1\mathrm{e}{-4}]$. Similarly to ImageNet, we use the best-performing checkpoint in the tests and use a ten-step Bayesian Optimization scheme to select performant hyperparameters.
@@ I.1 Log Probability Proper Scoring Rule for the Predictive
First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [38].
@@ I.2 Expected Calibration Error
To set up the required quantities for the Expected Calibration Error (ECE) metric [40], we follow the steps below, based on [38].
@@
The Expected Calibration Error (ECE) metric [40] is then defined as
@@ I.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
One can show that this is indeed a strictly proper scoring rule [38].
@@ I.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
Spectral normalised Gaussian processes (SNGP) [30] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty. For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
@@
The method applies spectral normalization to the hidden weights in each layer using a power iteration scheme with a single iteration per batch to obtain the largest singular value. Liu et al. [30] claim this helps with input distance awareness.
@@
Heteroscedastic classifiers (HET) [7] construct a Gaussian distribution in the logit space to model per-input uncertainties:
@@
The Laplace approximation [11] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian. In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularised loss over the training set $\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
@@
This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [48].
For deep neural networks, the Hessian matrix is often replaced with the Generalised Gauss-Newton (GGN) matrix. The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [36] or low-rank approximations.
@@
We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [48].
@@
Table˜K.3 shows results on ViT Little [14] backbones from the timm [52] library on the ImageNet validation set. As in the main paper, the hyperparameters are optimised using a ten-step Bayesian hyperparameter optimization scheme of Weights and Biases.
@@
We first evaluate calibration using the log probability scoring rule [17] and the Expected Calibration Error (ECE) metric [40]. Fig.˜3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
@@
Closed-form predictives do not sacrifice accuracy. Fig.˜K.3 evidences this claim on CIFAR-10: our closed-form predictives either outperform or are on par with Softmax predictives. The most accurate method is Sigmoid HET. These results support the findings of Wightman et al. [51] that showcase desirable training dynamics of the class-wise cross-entropy loss.
@@
Table˜K.8 provides results for two evidential deep learning methods, EDL [44] and PostNet [6]. (We adopt the naming convention of [mucsányi2024benchmarking].) We make the comparison because these methods also yield closed-form predictives and second-order quantities of interest, as they build Dirichlet distributions whose advantages we discuss in Section˜4.1. Following our experimental setup (Section 6), we optimise the hyperparameters of both methods using a ten-step Bayesian Optimization scheme in Weights & Biases. Compared with the results presented in the main text, our approach outperforms the two EDL methods across all metrics, significantly so on NLL and OOD AUROC. The results on the latter metric indicate that our moment-matched Dirichlet distributions capture second-order epistemic information more effectively than the loss-based EDL methods, which aligns with the findings of Bengs et al. [4].